Nick traveled out of the country for a month. He purchased a plan from his cell phone company that would allow him to send text messages to friends and family back home.​ Nick's telephone company charges $ 0.30

for a regular text message and $ 0.50

for a multimedia text messages. This month Nick was billed $ 89

and he sent total of 238

messages. How many regular text messages and multimedia text messages did he​ send?

Answer by Boreal(15235)

x=regular
238-x=multimedia
.30x+.50(238-x)=89
.30x-.50x+119=89
-.20x=-30
x=150 regular ($45)
238-x=88=multimedia ($44)
.